OpenStudy (anonymous):

7 1/5 - 2 3/10 = ? A. 5 3/4 B. 6 1/2 C. 6 D. 5 1/2

OpenStudy (ganpat):

7 1/5 = ( 7*5 + 1 )/ 5 = 36/5 find for other term and add them..